# The Machine Learning Control Method for Counterfactual $*{}^{*}$

### v1

@@ Abstract
Without a credible control group, the most widespread methodologies for estimating causal effects cannot be applied. To fill this gap, we propose the Machine Learning Control Method (MLCM), a new approach for causal panel analysis based on counterfactual forecasting with machine learning. The MLCM estimates policy-relevant causal parameters in short- and long-panel settings without relying on untreated units. We formalize identification in the potential outcomes framework and then provide estimation based on supervised machine learning algorithms. To illustrate the advantages of our estimator, we present simulation evidence and an empirical application on the impact of the COVID-19 crisis on educational inequality in Italy. We implement the proposed method in the companion R package MachineControl.
Keywords: potential outcomes framework, counterfactual forecasting, machine learning, short panel, panel cross-validation, education inequality
JEL-Codes: C13, C18, C53, I24
We are grateful to Guido Imbens and Fabrizia Mealli for thoughtful conversations and discussions. We also thank Andrea Albanese, Guglielmo Barone, Michele Battisti, Alessio D’Ignazio, Christina Gatmann, Anna Gottard, Giulio Grossi, Martin Huber, Michael Knaus, Joanna Kopinska, Giuseppe Maggio, Alessandra Mattei, Raffaele Mattera, Giovanni Mellace, Andrea Mercatanti, Samuel Nocito, Jason Poulos, Donato Romano, Jacques-François Thisse, Luca Tiberti, Giuseppe Ragusa and Giuliano Resce for valuable suggestions on earlier versions of this work. The paper has also benefited from many helpful comments and suggestions by audiences at the 2023 LISER International Workshop on “Machine Learning in Program Evaluation, High-dimensionality and Visualization Techniques”, the Oslo 2023 European Causal Inference Meeting, the AISRe, COMPIE, and SIE 2022 annual conferences and seminar participants at the Florence Center for Data Science and Sapienza University of Rome. We thank Gabriele Pinto and Federico Rucci for useful comments on the implementation.
@@ 1 Introduction
Today, the econometric toolbox for causal panel analysis features many alternative approaches for estimating causal effects, such as difference-in-differences (Card and Krueger,, 1994), the synthetic control method (Abadie et al.,, 2010), two-way (Angrist and Pischke,, 2009) and interactive (Bai,, 2009) fixed effects models, and matrix completion methods (Athey et al.,, 2021). However, all of these popular methodologies depend on a critical requirement: the availability of untreated units forming a credible control group, without which they cannot be applied. This poses a relevant econometric challenge in observational studies, as there are at least two relevant cases in which a control group does not exist: (i) the treatment simultaneously affects all units, as in the case of a large-scale shock or a nationwide program with universal participation (Duflo,, 2017) [^1] This case also encompasses situations where the treatment simultaneously affects most units, and the remaining ones are too few and too different to constitute a valid control group (even if using synthetic control methods).; (ii) only a subgroup of units gets treated, but the set of untreated units cannot form a valid control group due to violations of the no-interference assumption (Cox,, 1958). [^2] Spillovers are common in many economic settings where the observations are correlated in time, space, or both, and ignoring them can lead to very misleading inferences (Sobel,, 2006). In principle, spillovers can be modeled and accounted for at the cost of additional and often restrictive assumptions. However, in practice, most methods just postulate the absence of interference. Under these challenging but not uncommon circumstances, the use of standard causal panel data methods is precluded, leaving a gap in the econometric toolkit of empirical researchers.
@@
We fill this gap by introducing the Machine Learning Control Method (MLCM), a new estimator based on flexible counterfactual forecasting via machine learning (ML). The MLCM is a versatile technique that can leverage any available supervised ML algorithm and enables the estimation of several policy-relevant causal parameters—including individual, average, and conditional average treatment effects (CATEs)—in evaluation settings with short and long panels and no control group. The intuition behind the MLCM is straightforward: when you cannot rely on untreated units to build the counterfactual, you can forecast it. Specifically, use supervised ML techniques trained and tuned on the pre-treatment data to forecast the evolution of post-treatment outcomes in the absence of the treatment. Then, estimate treatment effects as the difference between observed and forecasted post-treatment outcomes.
@@
Our approach builds upon and bridges three different methodological currents: causal panel data methods, time series forecasting, and causal ML. From the literature on causal panel analysis, we take the conceptual and identification framework based on the potential outcomes model (Rubin,, 1974). Since earlier contributions (Ashenfelter and Card,, 1985; Card,, 1990) to its more recent developments (Arkhangelsky et al.,, 2021; Roth et al.,, 2023), this whole literature ultimately revolves around solving the “fundamental problem of causal inference” (Holland,, 1986) in an observational panel data setting. To this end, all existing methodologies leverage untreated units to estimate the counterfactual scenario, i.e., how the outcome of the treated group would have evolved in the absence of the treatment. After building a credible control group, the causal effect is recovered as the difference in outcomes between the two groups. Central to our work is also the perspective of causal inference as a missing data problem, where the key goal is the imputation of missing potential outcomes (Imbens and Rubin,, 2015). We contribute to this literature by introducing a novel estimator that imputes missing potential outcomes and, in turn, gauges causal effects without relying on untreated units.
@@
The time series literature features some forecasting approaches that researchers have employed for counterfactual building in scenarios without a control group. Some of them are straightforward: the mean method estimates the counterfactual “no-treatment scenario” of the dependent variable $\operatorname{Y}_{i}$ as the average of its pre-treatment values, while the naive method considers the last pre-treatment value of $\operatorname{Y}_{i}$ as the counterfactual (Hyndman and Athanasopoulos,, 2021). A more sophisticated forecasting approach is the interrupted time series (ITS) analysis, first formalized by Box and Tiao, (1975). ITS analysis generally fits regression models or autoregressive integrated moving average (ARIMA) models to the entire time series of observed data; this is done after postulating a structure on the intervention effect (e.g., constant level shift). By construction, ITS delivers a single average effect for the whole post-intervention period. More recent studies (Brodersen et al., (2015), Chernozhukov et al., (2021), Menchetti et al., (2023)) have formalized estimation and inference with time series forecasting techniques (such as ARIMA) within the potential outcomes framework. However, such methods are designed for cases with a single treated series, often impose restrictive functional forms, and can be implemented only when at least tens of pre-treatment periods are available. [^3] Closely related to these time series approaches, Botosaru et al., (2023) propose the use of simple techniques, such as polynomial regressions, to estimate treatment effects one period after the implementation of the treatment in panel settings without controls. We draw from this literature the idea of forecasting the counterfactual when no control group is available and the estimation approach based on temporal cross-validation. We add to it by proposing a flexible, ML-powered forecasting technique suitable for panel data and able to estimate separate effects for multiple post-intervention periods.
@@
In recent years, a new literature at the intersection between causal inference with panel data and ML has evolved. Causal trees and forests (Athey and Imbens, (2016), Wager and Athey, (2018)) constitute an adaptation of tree-based predictive models to the task of estimating heterogeneous treatment effects, and can be applied to both randomized and observational data. Although they are not specifically designed for panel settings, they can also be employed with panel data (e.g., Athey et al., (2023); Britto et al., (2022)). Artificial control methods (Carvalho et al., (2018), Masini and Medeiros, (2021)) and the synthetic learner proposed by Viviano and Bradic, (2023) harness supervised ML algorithms to predict counterfactuals and assess treatment effects in a panel setting with a single treated unit, many untreated peers, and a long pre-intervention window. Nevertheless, in order to build a counterfactual scenario, all these causal ML techniques need a set of units assumed to be completely unaffected by the treatment. We borrow from this literature the idea of harnessing the power of ML in the service of causal inference rather than for pure prediction, exploiting the fact that counterfactual building is ultimately a predictive task (Varian,, 2016). We extend it by proposing a ML estimator that: i) does not depend on the availability of untreated units or on the no-interference assumption; ii) is designed for situations with many treated units; iii) also works with a few pre-treatment periods.
@@
After discussing identification in the potential outcomes framework, we propose an estimation strategy based on a horse-race competition between several different ML algorithms, panel cross-validation (CV) for model selection, and block-bootstrapping for inference. The MLCM comes with a set of diagnostic, performance, and placebo tests and is characterized by a high level of generality: it delivers individual treatment effects that can either be the object of interest or can later be aggregated into several policy-relevant causal estimands. Since our approach allows for arbitrary and unrestricted treatment effect heterogeneity, we also propose an automated search for heterogeneity based on an easy-to-interpret regression tree. To showcase the potential of the MLCM, we present a simulation study and an empirical application in which we investigate the effects of the COVID-19 pandemic on educational inequality in Italian Local Labor Markets (LLMs). We find that the pandemic led to a generalized drop in students’ performance, which is particularly pronounced in LLMs that before COVID-19 were characterized by higher levels of unemployment and inequality and lower educational attainment.
@@
The possibility of moving beyond the reliance on untreated units paves the way to evaluating many potential treatments—such as universal policies, large-scale shocks, and programs engendering interactions between units—that, due to the lack of a valid comparison group, have so far been underexplored in empirical studies. To answer these causal questions, interested researchers can implement the MLCM on real-world datasets using the companion R package MachineControl. [^4] The latest version of the package is available on GitHub at this link.
@@
The rest of this paper is organized as follows. Section 2 outlines the causal framework by defining the identification assumptions, the causal estimands, and the estimators. Section 3 describes the implementation process of the MLCM and presents the simulation study. Section 4 illustrates the empirical application. Section 5 concludes.
@@ 2 The causal framework
In this section, we present the causal framework for an observational short-panel setup where the intervention is a simultaneous policy change or shock that affects directly or indirectly the entire statistical population under scrutiny. [^5] We take the benchmark case of a balanced panel, but the MLCM accommodates also the use of unbalanced panels. In such a setting, the estimation of causal effects must rely on estimators specifically designed for scenarios without a control group. We start by defining the causal assumptions and the corresponding causal estimands within the potential outcomes framework. We then provide the proof that such estimands can be identified under those assumptions. Finally, we specify the modeling assumptions and introduce novel estimators for scenarios without untreated units. To make the notation (and the method) as general as possible, the definitions given in this section assume the presence of some contemporaneous covariates unaffected by the intervention. Including these covariates in the MLCM can potentially mitigate certain assumptions. However, as we will discuss later, caution must be exercised in their use and selection.
2.1 Causal assumptions
We need three assumptions to define, identify, and estimate the causal estimands of interest.
@@ 2.1 Causal assumptions
Denote with $\operatorname{Y}_{i,t}$ the outcome of unit $i$ at time $t$, and let $\operatorname{W}_{i,t}\in\{0,1\}$ be a random variable describing the treatment assignment of unit $i=1,\dots,N$ at time $t=1,\dots,t_{0},\dots,T$, where $1$ indicates the treatment, 00 indicates control, and $t_{0}$ denotes the intervention date. [^6] The word “treatment” is commonly used in the context of randomized controlled trials. As we are dealing with an observational study, we use the terms “treatment”, “intervention”, “policy”, and “shock” interchangeably. As we are focusing on a single and simultaneous intervention, we can then write $\operatorname{W}_{i,t}=0$ for all $i$ and $t\leq t_{0}$, $\operatorname{W}_{i,t}=1$ for all $i$ and $t>t_{0}$, therefore, $t_{0}+1$ is the first post-intervention period. In other words, the units become treated at the same time and then stay treated. Under the potential outcomes framework, for each unit we can define a potential outcome under $\operatorname{W}_{i,t}=0$ and a different potential outcome under $\operatorname{W}_{i,t}=1$.
@@
The first assumption establishes the crucial link between potential outcomes and treatment assignment. We rely on a less stringent version of the Stable Unit Treatment Value Assumption (SUTVA) (Rubin,, 1974; Imbens and Rubin,, 2015). Specifically, we substantially relax SUTVA, retaining only its second part: the treatment is the same for all the units. This is an a priori assumption that the potential outcomes do not depend on different treatment intensities or the mechanism used to assign the treatment.
Assumption 1
@@
While we maintain this no-multiple-versions-of-treatment assumption, we drop the first part of SUTVA, namely, the no-interference assumption (Cox,, 1958). [^7] Following Imbens and Rubin, (2015), we consider the case with general equilibrium effects as a scenario in which there are widespread violations of the no-interference assumption. Hence, general equilibrium effects are an example of substantial interaction between units, rather than a separate category. We consider this as one of the main advantages of our approach because, while it has long been known that violations and failures of the no-interference assumption can lead to misleading inferences in many social science settings (Sobel,, 2006), this strong assumption is rarely questioned or tested in practice (Chiu et al.,, 2023). This is a key departure from most evaluation methods and, therefore, it is important to delve into its implications. SUTVA-related interference among units can be of two types: (1) from treated to control units, and (2) among treated units themselves. [^8] Even if the treatment is the same for all units, residual spillover effects may arise due to interference among the treated units. Consider, for example, a study investigating excess mortality from COVID-19. While the pandemic affects all areas, those with fewer intensive care beds may need to send patients to neighboring hospitals, prompting an additional (indirect) rise in mortality rates there also. Spillovers due to individuals’ characteristics within the same treatment group are discussed in Ogburn and VanderWeele, (2014). First, we completely circumvent pitfalls regarding Type-1 interference, since our counterfactual scenario is generated without relying on untreated units. Second, we do not postulate the lack of interference among treated units and allow for any possible Type-2 interference. We avoid relying on the no-interference-among-treated assumption because we are aware of the likely presence of interference across both the temporal and cross-sectional dimensions in social science applications (Xu,, 2023). Our focus is on estimating the total effect of the treatment, rather than on disentangling its direct and indirect effects. The total effect encompasses both the direct effect on a unit from the treatment it receives and the indirect effects arising from the spillover and general equilibrium effects. We claim that in real-world scenarios with likely interaction between units, focusing on the total effect of the treatment is a sensible choice as it captures the actual impact on each unit under the realized treatment assignment mechanism. [^9] In the rare cases in which there will be no Type-2 interference, then the MLCM will retrieve the direct effect of the treatment on each unit, which will coincide with the unit-specific total effect. Under Assumption 1 we can then write $\operatorname{Y}_{i,t}(w_{i,t})$ to indicate the potential outcomes, where the lower case $w_{i,t}$ denotes a realization of $\operatorname{W}_{i,t}$.
@@
The next assumption rules out the possibility of anticipatory effects of the intervention on both the outcome and the covariates (Abadie et al.,, 2010; Borusyak et al.,, 2023).
Assumption 2
@@
Let $\mathbf{X}_{i,t}=(\operatorname{X}_{i,t}^{(1)},\dots,\operatorname{X}_{i,t}^{(m)})$ be an $m$-dimensional vector of covariates that are predictive of the outcome $i$ at time $t\leq t_{0}$; both the covariates and the potential outcomes are unaffected by the policy for all $t\leq t_{0}$, i.e., $\operatorname{Y}_{i,t}(1)=\operatorname{Y}_{i,t}(0)$ and $\mathbf{X}_{i,t}(1)=\mathbf{X}_{i,t}(0)$. Some of the covariates are also unaffected by the policy in the post-intervention period, i.e., for all $t>t_{0}$, $\mathbf{X}_{i,t}(1)=\mathbf{X}_{i,t}(0)$.

### v2

# Causal inference and policy evaluation without a control group

@@ Abstract
Without a control group, the most widespread methodologies for estimating causal effects cannot be applied. To fill this gap, we propose the Machine Learning Control Method, a new approach for causal panel analysis that estimates causal parameters without relying on untreated units. We formalize identification within the potential outcomes framework and then provide estimation based on machine learning algorithms. To illustrate the practical relevance of our method, we present simulation evidence, a replication study, and an empirical application on the impact of the COVID-19 crisis on educational inequality. We implement the proposed approach in the companion R package MachineControl.
Keywords: potential outcomes framework, counterfactual forecasting, machine learning, short panels, panel cross-validation, educational inequality.
JEL-Codes: C18, C21, C53, I24.
——————— Department of Social Sciences and Economics, Sapienza University of Rome, Rome, Italy (IT). Email at: augusto.cerqua@uniroma1.it; marco.letta@uniroma1.it
Department of Statistics, Computer Science and Applications, University of Florence, Florence, Italy (IT). Email at: fiammetta.menchetti@unifi.it
*We are grateful to Guido Imbens and Fabrizia Mealli for thoughtful conversations and discussions. We also thank Andrea Albanese, Guglielmo Barone, Michele Battisti, Iavor Bojinov, Fabrizio Cipollini, Alessio D’Ignazio, Christina Gatmann, Anna Gottard, Giulio Grossi, Martin Huber, Michael Knaus, Joanna Kopinska, Giuseppe Maggio, Alessandra Mattei, Raffaele Mattera, Giovanni Mellace, Andrea Mercatanti, Samuel Nocito, Gabriele Pinto, Jason Poulos, Donato Romano, Federico Rucci, Jacques-François Thisse, Luca Tiberti, Giuseppe Ragusa, Giuliano Resce, and Bas van der Klaauw for valuable suggestions on earlier versions of this work. The paper has benefited from helpful comments and suggestions by audiences at many conferences and seminars.
@@ 1 Introduction
Today, the econometric toolbox for causal panel analysis features many alternative approaches for estimating causal effects, such as difference-in-differences (Card and Krueger, 1994), the synthetic control method (Abadie et al., 2010), two-way (Angrist and Pischke, 2009) and interactive (Bai, 2009) fixed effects models, and matrix completion methods (Athey et al., 2021). However, all of these popular methodologies depend on a critical requirement: the availability of untreated units forming a credible control group, without which they cannot be applied. This poses a significant empirical challenge in observational studies, as there are at least two relevant cases in which a suitable control group does not exist: (i) when the treatment simultaneously affects all units, such as in the context of a large-scale shock or a nationwide program with universal participation (Duflo, 2017); (ii) when only a subgroup of units receives the treatment, but the set of untreated units cannot form a valid control group due to violations of the no-interference assumption (Cox, 1958). [^1] Spillovers are common in many empirical settings where the observations are correlated in time, space, or both, and ignoring them can lead to very misleading inferences (Sobel, 2006). In principle, spillovers can be modeled and accounted for at the cost of additional and often restrictive assumptions. However, in practice, most methods just postulate the absence of interference. Under these challenging but not uncommon circumstances, the use of standard causal panel data methods is precluded, leaving a gap in the econometric toolkit of empirical researchers.
@@
We fill this gap by introducing the Machine Learning Control Method (MLCM), a new estimator based on flexible counterfactual forecasting via machine learning (ML). The MLCM is a versatile technique that can leverage any available supervised ML algorithm and enables the identification and estimation of several policy-relevant causal parameters—including individual, average, and conditional average treatment effects (CATEs)—in practically relevant environments without a control group. It is flexible in terms of data structure, as it can be applied across a wide variety of panel settings, including very short panels and contexts with staggered adoption. The intuition behind the MLCM is straightforward: when you cannot rely on untreated units to build the counterfactual, you can forecast it. Specifically, use supervised ML techniques trained and tuned on the pre-treatment data to forecast the evolution of post-treatment outcomes in the absence of the treatment. Then, estimate treatment effects as the difference between observed and forecasted post-treatment outcomes.
@@
Our approach builds upon and bridges three different methodological currents: causal panel data methods, time series forecasting, and causal ML. From the literature on causal panel analysis, we adopt the conceptual and identification framework based on the potential outcomes model (Rubin, 1974). Since earlier contributions (Ashenfelter and Card, 1985; Card, 1990) to its more recent developments (Arkhangelsky et al., 2021; Roth et al., 2023), this entire literature revolves around addressing the “fundamental problem of causal inference” (Holland, 1986) in an observational panel data setting where, for each unit, it is impossible to observe both potential outcomes at once. Therefore, causal inference can be viewed as a missing data problem, with the key goal being the imputation of missing potential outcomes for the treated units (Imbens and Rubin, 2015). Most widely used causal panel data methods achieve this imputation by leveraging untreated units to construct a control group. Our key contribution to this literature is to establish identification conditions in the absence of a control group and to introduce a novel estimator of causal effects that imputes missing potential outcomes without relying on untreated units.
@@
The time series literature features some forecasting approaches that researchers have employed for counterfactual building in scenarios without a control group. [^2] A rich tradition also exists for forecasting with panel data (e.g., Arellano and Bonhomme (2011); Baltagi (2008); Liu et al. (2020)). However, our interest lies in counterfactual forecasting. In this regard, the Rubin causal framework is absent from this tradition, which focuses on different estimands and non-causal settings compared to the recent causal panel data literature (Arkhangelsky and Imbens, 2024). Some of them are straightforward: the mean method estimates the counterfactual “no-treatment scenario” of the dependent variable $\operatorname{Y}_{i}$ as the average of its pre-treatment values, while the naive method considers the last pre-treatment value of $\operatorname{Y}_{i}$ as the counterfactual (Hyndman and Athanasopoulos, 2021). A more sophisticated forecasting approach is the interrupted time series (ITS) analysis, first introduced by Box and Tiao (1975). ITS analysis generally fits regression models or autoregressive integrated moving average (ARIMA) models to the entire time series of observed data; this is done after postulating a structure on the intervention effect (e.g., constant level shift). By construction, ITS delivers a single average effect for the whole post-intervention period. More recent studies (Brodersen et al., 2015; Chernozhukov et al., 2021; Menchetti et al., 2023; Rambachan and Shephard, 2021) consider estimation and inference with other time series techniques (such as Bayesian Structural Time Series, ARIMA models, or impulse response functions) in time series settings where no direct control group may be available. From a panel data perspective, these methods have significant limitations. First, they are designed for cases with a single treated series. Second, they often impose restrictive assumptions and functional forms. Third, they can be implemented only when many pre-treatment periods are available. However, in the traditional panel data literature, the number of units is much larger than the number of time periods (Arkhangelsky and Imbens, 2024), and the time dimension is typically substantially smaller than in time series settings, rendering these methods mostly inapplicable. We draw from this literature the idea of forecasting counterfactuals by exploiting the temporal information. We go beyond it by proposing a more flexible, ML-powered method explicitly designed for counterfactual forecasting with panel data.
@@
In recent years, a new methodological literature at the intersection between causal inference with panel data and ML has evolved. The matrix completion methods developed by Athey et al. (2021) use nuclear norm regularization and observed control outcomes to impute missing elements in the counterfactual matrix of untreated unit/period combinations. Semenova et al. (2023) propose an approach for estimating CATEs characterized by high-dimensional parameters in both homogeneous cross-sectional and unit-heterogeneous dynamic panel data settings, leveraging ML techniques for improved inference. Artificial control methods (Carvalho et al., 2018; Masini and Medeiros, 2021) and synthetic learners (Viviano and Bradic, 2023) harness supervised ML algorithms to predict counterfactuals and assess treatment effects in a panel setting with a single treated unit, many untreated units, and a long pre-intervention window. All these methods need a set of units assumed to be completely unaffected by the treatment to build a counterfactual scenario. Therefore, none of these causal ML techniques can be applied in the challenging econometric setting we focus on. We share with this literature the idea of harnessing the power of ML in the service of causal inference rather than for pure prediction, exploiting the fact that counterfactual building is ultimately a predictive task (Varian, 2016). We extend it by developing a new causal ML estimator that estimates individual treatment effects without depending on the no-interference assumption or leveraging unaffected units.
@@
We begin by thoroughly discussing and formalizing identification without a control group in the potential outcomes framework. We then propose an adaptive estimator based on a horse-race competition between several different ML algorithms, a novel cross-validation (CV) procedure for model assessment and selection tailored for panel data, and block-bootstrapping for inference. The MLCM comes with a set of diagnostic and placebo tests and is characterized by a high level of generality: it delivers individual treatment effects that can either be the object of interest or aggregated into several policy-relevant causal estimands. Since our approach allows for unrestricted treatment effect heterogeneity, we also propose an automated search for heterogeneity based on an easy-to-interpret regression tree.
To showcase the potential of the MLCM, we present an extensive simulation study, a replication of the minimum wage application in Callaway and Sant’Anna (2021) without using control units, and an empirical application in which we investigate the effects of the COVID-19 pandemic on educational inequality in Italian Local Labor Markets (LLMs). We find that the pandemic led to a generalized drop in students’ performance, which is particularly pronounced in LLMs that before COVID-19 were characterized by higher levels of unemployment and inequality and lower educational attainment.
At the time of writing, we are aware of only one other method for assessing causality in panel settings without using control units. [^3] There are also some empirical studies in energy economics that, without proposing a new formal method, have applied counterfactual prediction using ML algorithms on very high-frequency (hourly) electricity data (Abrell et al., 2022; Jarvis et al., 2022; Prest et al., 2023). In independent work subsequent to ours, Botosaru et al. (2023) propose estimating average treatment effects in the absence of a control group by regressing pre-treatment outcomes on known basis functions of time. Compared to the MLCM, this method relies on different assumptions, such as absence of interference and the validity of a central limit theorem, and imposes functional form restrictions. Overall, we view the two methods as complementary, depending on the assumptions one is willing to make and on the characteristics of the available data. [^4] For instance, the estimator by Botosaru et al. (2023) can be employed in data-scarce environments where only outcome data are available, or when the researcher is comfortable relying on parametric assumptions. In contrast, the MLCM is more flexible and better suited for data-rich or high-dimensional settings where information on many covariates is available. This context is particularly relevant when the primary goal is to uncover heterogeneity in treatment effects associated with differences in observed characteristics.
@@
The possibility of moving beyond the reliance on untreated units paves the way to evaluating many potential treatments—such as universal policies, large-scale shocks, and programs engendering interactions between units—that, due to the lack of a valid comparison group, have so far been under-explored in empirical studies. To answer these causal questions, interested researchers can implement the MLCM on real-world datasets using the companion R package MachineControl. [^5] The latest version of the package is available on GitHub at this link.
@@
The rest of this paper is organized as follows. Section 2 outlines the causal framework by defining the identification assumptions, the causal estimands, and the estimators. Section 3 describes the implementation process of the MLCM and presents the simulation study. Section 4 illustrates the empirical application, while Section 5 concludes.
@@ 2 The causal framework
In this section, we present the causal framework for an observational short-panel setup where the intervention is a simultaneous policy change or shock that affects directly or indirectly the entire statistical population under scrutiny. [^6] To use language from Arkhangelsky and Imbens (2024), in this benchmark case, we consider a thin panel matrix, with a large number of cross-sectional units and a small number of time periods. However, the framework easily accommodates the use of long panels and fat and square matrices. In addition, the MLCM can be applied to both balanced and unbalanced panels, as well as in block-assignment and staggered adoption settings where one cannot credibly rely on the plausibility of the no-interference assumption (i.e., outcomes of never-treated or not-yet-treated units are contaminated by spillovers). We start by defining the assumptions and the corresponding causal estimands within the potential outcomes framework. We then provide the proof that such estimands can be identified under those assumptions. Finally, we introduce novel estimators for scenarios without a suitable control group. To make the notation (and the method) as general as possible, the definitions given in this section assume the presence of some contemporaneous covariates unaffected by the intervention. Including these covariates in the MLCM can potentially mitigate certain assumptions. However, as we will discuss later, caution must be exercised in their use and selection.
2.1 Assumptions
@@
Denote with $\operatorname{Y}_{i,t}$ the outcome of unit $i$ at time $t$, and let $\operatorname{W}_{i,t}\in\{0,1\}$ be a random variable describing the treatment assignment of unit $i=1,\dots,N$ at time $t=1,\dots,t_{0},\dots,T$, where $1$ indicates the treatment, 00 indicates control, and $t_{0}$ denotes the intervention date. [^7] The word “treatment” is commonly used in the context of randomized controlled trials. As we are dealing with an observational study, we use the terms “treatment”, “intervention”, “policy”, and “shock” interchangeably. As we are focusing on a single and simultaneous intervention, we can then write $\operatorname{W}_{i,t}=0$ for all $i$ and $t\leq t_{0}$, $\operatorname{W}_{i,t}=1$ for all $i$ and $t>t_{0}$, therefore, $t_{0}+1$ is the first post-intervention period. In other words, all units are treated simultaneously and remain treated thereafter. Under the potential outcomes framework, for each unit, we can define a potential outcome under $\operatorname{W}_{i,t}=0$ and a different potential outcome under $\operatorname{W}_{i,t}=1$.
We now introduce the assumptions that are needed to define, identify, and estimate causal effects in an observational panel setting without controls, offering novel theoretical insights such as identification conditions for non-linear, multi-step-ahead causal effects. We emphasize that such assumptions are not particularly restrictive, as the MLCM is designed to be highly flexible. Our method, whose implementation is detailed in Section 3, starts by running several ML algorithms on pre-treatment data and then selects the one producing the most accurate forecasts. Given that it is not possible to know in advance which ML method will be chosen, the notation is kept as general as possible. [^8] We believe this is an advantage over most existing approaches in the causal ML literature: while Carvalho et al. (2018) and Masini and Medeiros (2021) focus on LASSO and Wager and Athey (2018) adopt tree-based methods, the MLCM is flexible and can easily adapt to different data-generating processes. A notable exception in this context is the synthetic learner proposed by Viviano and Bradic (2023). In contrast to their technique, our method is based on a horse-race competition between several alternative ML algorithms, whereas the synthetic learner relies on an ensemble procedure that combines various estimators.
@@
The first assumption contributes to define potential outcomes and establishes the crucial link between them and the treatment assignment. We rely on a less stringent version of the Stable Unit Treatment Value Assumption (SUTVA) (Rubin, 1974; Imbens and Rubin, 2015), retaining only its second part: the treatment is the same for all the units. This is an a priori assumption that the potential outcomes do not depend on different treatment intensities or the mechanism used to assign the treatment.
Assumption 1 (Weak SUTVA).
@@
While we maintain this no-multiple-versions-of-treatment assumption, we drop the first part of SUTVA, namely, the no-interference assumption (Cox, 1958). [^9] Following Imbens and Rubin (2015), we consider the case with general equilibrium effects as a scenario in which there are widespread violations of the no-interference assumption. We consider this as one of the main advantages of our approach because, while it has long been known that violations and failures of the no-interference assumption can lead to misleading inferences in many social science settings (Sobel, 2006), this strong assumption is rarely questioned or tested in practice (Chiu et al., 2023). This is a key departure from most evaluation methods and it is important to delve into its implications. SUTVA-related interference among units can be of two types: (1) from treated to control units, and (2) among treated units themselves. [^10] Even if the treatment is the same for all units, residual spillover effects may arise due to interference among the treated units. Consider, for example, a study investigating excess mortality from COVID-19. While the pandemic affects all areas, those with fewer intensive care beds may need to send patients to neighboring hospitals, prompting an additional (indirect) rise in mortality rates there also. Spillovers due to individuals’ characteristics within the same treatment group are discussed in Ogburn and VanderWeele (2014). First, we completely circumvent pitfalls regarding Type-1 interference, since our counterfactual scenario is generated without relying on untreated units. Second, we do not postulate the lack of interference among treated units and allow for any possible Type-2 interference. We avoid relying on the no-interference-among-treated assumption because of the likely presence of interference across both the temporal and cross-sectional dimensions in social science applications (Xu, 2023). Consequently, our focus is on estimating the total effect of the treatment, which encompasses both the direct effect on a unit from the treatment it receives and the indirect effects arising from the spillover and general equilibrium effects. We claim that in real-world scenarios with likely interaction between units, focusing on the total effect of the treatment is a sensible choice as it captures the actual impact on each unit under the realized treatment assignment mechanism. Under Assumption 1, we can index the potential outcomes to the common treatment path: $\operatorname{Y}_{i,t}(1)$ indicates the potential outcome under the treatment, and $\operatorname{Y}_{i,t}(0)$ is the counterfactual outcome.
The next assumption contributes to the definition of causal effects.
Assumption 2 (Additivity).
We assume that the intervention produces an additive effect on the potential outcomes, i.e.,
$\operatorname{Y}_{i,t}(1)=\operatorname{Y}_{i,t}(0)+\Delta_{i,t}.$ (1)
This is not a restrictive assumption, since it holds after a suitable transformation of the data. For example, we can start from a multiplicative effect and then apply a logarithmic transformation to find an additive structure. Since inference is conducted using block-bootstrap, if it is necessary to express the effect in its original form, the inverse transformation can be applied to the bootstrap distribution, so as to recover the original multiplicative effect and its confidence interval. In conjunction with Assumption 4 (see below), this assumption implies that we effectively adopt a Partially Linear Model, akin to seminal causal machine learning methods (Chernozhukov et al., 2018; Wager and Athey, 2018). The Partially Linear Model strikes a balance between structure and flexibility: the causal-effect component of the model remains simple and interpretable, while the untreated potential outcome can exhibit almost arbitrary complexity (Chernozhukov et al., 2024).
@@
Assumption 3 rules out the possibility of anticipatory effects of the intervention on both the outcome and the covariates (Abadie et al., 2010; Borusyak et al., 2023).
Assumption 3 (Absence of anticipation).
@@
Let $\mathbf{X}_{i,t}=(\operatorname{X}_{i,t}^{(1)},\dots,\operatorname{X}_{i,t}^{(m)})$ be an $m$-dimensional vector of covariates that are predictive of the outcome $i$ at time $t\leq t_{0}$. We assume: i) absence of anticipatory effects of the intervention on the covariates and the potential outcomes, i.e., $\operatorname{Y}_{i,t}(1)=\operatorname{Y}_{i,t}(0)$ and $\mathbf{X}_{i,t}(1)=\mathbf{X}_{i,t}(0)$ for all $t\leq t_{0}$; ii) future covariates do not affect current potential outcomes; iii) (optional) some of the covariates remain unaffected by the policy in the post-intervention period (post-treatment exogeneity of the covariates), i.e., for all $t>t_{0}$, $\mathbf{X}_{i,t}(1)=\mathbf{X}_{i,t}(0)$.